7. Two batteries and three resistors are connected as shown below. 5.0 ? 6.0 2 4.0 ? How much current flows through each resistor when the switch is a. open? b. closed?
Answer #1

7.

(a)

6.0 Ω 4.0 Ω

9-6-6I-4I=0

I=0.3\: A (Same for both 6 ohms and 4 ohms resistor)

Zero current through 5 ohms resistor.

(b)

9.0 V 5.0 Ω 6.0 Ω 4.0 Ω

9-6+6I_2-4I_1=0

4I_1-6I_2=3.....................(I)

9-5(I_1+I_2)-4I_1=0

9I_1+5I_2=9.......................(II)

I_1=0.93\: A,I_2 = 0.12\: A

I_1+I_2= 1.05\: A